Three flasks, each containing a gas sample, are set up as shown below. Assuming no temperature change, determine the final pressure inside the system after all stopcocks are opened. Assume that the connecting tube has negligble volume.

A = CH4(g) V = 3.69 L P = 0.836 atm

B = H2(g) V = 2.76 L P = 1.10 atm

C = CO2(g) V = 5.09 L P = 2.51 atm

Final pressure = ______ atm.
1
Expert's answer
2019-03-18T16:24:41-0400

Total volume V = 3.69 + 2.76 + 5.09 = 11.54 (L)

p(CH4) = 0.836 * (3.69 / 11.54) = 0.267 (atm)

p(H2) = 1.10 * (2.76 / 11.54) = 0.263 (atm)

p(CO2) = 2.51 * (5.09 / 11.54) = 1.11 (atm)

p(final) = 0.267 + 0.263 + 1.11 = 1.64 (atm)

Answer: p = 1.64 atm